OpenStudy (anonymous):

This question is wanting me to provide what should be added to both sides of the equation x^2 +8x=-9

OpenStudy (anonymous):

just add the 9

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Im sorry, I just am no good at this

OpenStudy (anonymous):

no x^2 +8x=-9 x^2 +8x + 9 =-9 +9 x^2 +8x + 9 = 0 are you trying to graph it?

OpenStudy (anonymous):

no i just need to know what to add to each side of the equation

OpenStudy (anonymous):

yup thats it then, just add 9 to both sides and you get \[x^2 +8x + 9 = 0\]

OpenStudy (anonymous):

k Thank You, Thank Goodness I only have one more. I have been working on these for 5 hours now. Thanks again